# Entanglement suppression and quantum scars in a three-oscillator gravitational analogue

### v1

@@ Abstract
We study quantum correlations in a tripartite system of coupled harmonic oscillators, modeling mediator-induced interactions akin to those in quantum gravity proposals. Surprisingly, despite the system’s integrability, we identify regimes where bipartite entanglement between distant oscillators vanishes even with non-zero quantum couplings. In the Heavy Mediator Regime, entanglement is restricted to sharply localized islands in parameter space, surrounded by extended regions of suppression, analogous to quantum scars in non-integrable systems. Fidelity analysis reveal that these low-entanglement states exhibit dynamical stability and spectral signatures reminiscent of quantum scars in non-integrable systems, suggesting a continuous-variable analogue of scarring tied to hidden phase-space symmetries. The Light Mediator Regime, by contrast, displays smooth entanglement generation. These results challenge entanglement-based tests of quantum gravity: the absence of entanglement need not imply classical mediation but may instead reflect a mediator dynamically constrained to a quantum subspace—akin to gravitational memory or decoherence-free subspaces. Our findings emphasize the necessity of probing the mediator’s dynamical regime to unambiguously diagnose quantum gravity.
@@ I Introduction
The Bose-Marletto-Vedral (BMV) proposal [1, 2] has emerged as a promising avenue to experimentally probe the quantum nature of gravity. It proposes that two mesoscopic masses, each in a spatial superposition, would become entangled through their gravitational interaction if gravity is quantum, a signature absent in a classical gravitational field [3, 4, 5]. While conceptually elegant, the BMV protocol faces formidable experimental challenges, such as maintaining coherence and detecting extremely weak gravitational entanglement [6, 7, 8, 9, 10, 11]. More fundamentally, theoretical ambiguities persist, with studies questioning whether entanglement is an unambiguous signature of quantum gravity, as classical systems can mimic statistical correlations or Bell-local correlations under certain conditions, [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. These concerns highlight the need for robust observables to discriminate between classical and quantum gravitational interactions.
In this work, we present a simplified yet insightful model of tripartite coupled harmonic oscillators (CHOs) to investigate the limitations of using entanglement as a definitive witness for quantumness of interactions, particularly in the context of BMV proposal. We demonstrate that for specific parameter regimes, this integrable system exhibits vanishing logarithmic negativity ($\mathcal{E}_{\mathcal{N}}$[24] (and quantum discord ($\mathcal{D}$[25]), despite the presence of non-zero interactions and quantum states. This behavior stems from the delicate balance of coupling strengths and oscillator frequencies, showing that vanishing entanglement or discord does not preclude quantum nature of the underlying interactions or the non-classicality of the system’s state. Our model offers a tractable framework to identify potential pitfalls in entanglement-based tests of quantum gravity.
The phenomenon of vanishing entanglement in our integrable three CHOs model bears a sriking resemblance to quantum many-body scar-states in certain non-integrable systems [26, 27, 28]. Quantum scars are specific eigenstates that defy the eigenstate thermalization hypothesis [29], exhibiting unexpectedly low entanglement and persisting quantum coherences in otherwise chaotic systems. While scars arise due to fine-tuned parameters or symmetries in non-integrable models [30], our integrable system demonstrates that such low-entanglement states can also emerge from a small set of Gaussian states in the Hilbert space. Further fidelity ($\mathcal{F}$[31] of the evolved state with the initial state remains remarkably close to unity over long times for such states  [^1] While traditionally underexplored in the literature, recent studies have shown an interest in this aspect [55].. This coexistence of dynamically localized states points towards selective resonance effects or subtle hidden symmetries. Such mechanisms could complicate entanglement-based tests of quantum gravity.
Our results have implications extending beyond the BMV proposal, underscoring the necessity of developing alternative observables and refined criteria to discriminate between classical and quantum interactions. Crucially, we demonstrate that entanglement alone cannot serve as a definitive marker of quantumness — particularly in complex systems where quantum correlations may vanish despite non-trivial dynamics. Our analysis suggests that robust tests of quantum gravity must include a broader suite of quantum signatures, including non-classical correlations beyond entanglement and interaction dynamics. This work serves as a cautionary note against overreliance on simplistic quantum witnesses, highlighting the need for rigorous theoretical frameworks to interpret experimental tests of quantum gravity.
@@ II The model and set up
To explore these questions, we analyze a simple tripartite system of quantum CHOs, designed to mimic aspects of the BMV setup as shown in Fig. (1). We investigate the evolution of $\mathcal{E}_{\mathcal{N}}$ [24] and $\mathcal{D}$ [25], as functions of system parameters and initial conditions. The system consists of three CHOs, labeled A, B, and C. Oscillator B acts as a local mediator, connecting A and C, which do not interact directly. For simplicity, oscillators A and C are assumed identical.
The Hamiltonian incorporates three types of interactions: self-squeezing, parametric squeezing, and mode mixing. The interaction between the mediator and the other oscillators resembles those in the Caldeira-Leggett model, coupling subsystems via terms like $x_{A}x_{B}$ or $x_{B}x_{C}$ [33]. These inter-oscillator couplings decompose into two qualitatively distinct components: mode mixing, which conserves particle number (e.g., $a^{\dagger}b+ab^{\dagger}$), and parametric squeezing, which does not (e.g., $ab+a^{\dagger}b^{\dagger}$[34]. This decomposition facilitates non-trivial energy redistribution and building of correlations across subsystems. For instance, correlated two-mode processes can lead to simultaneous excitation loss from modes A and B, while mode C gains energy through indirect coupling. Although the total energy of the system is conserved, these interaction channels enable rich dynamics of information and entanglement flow among the oscillators.
We include self-squeezing terms in the Hamiltonian, similar to those observed in quantum optics for light propagating through non-linear media [35, 36]. Their inclusion is motivated by the non-linear nature of gravity, where even in the linear limit, phenomena like gravitational memory exist [37, 38, 39]. Thus, considering self-squeezing is crucial for a model aiming to shed light on gravitational interactions. This framework allows us to explore scenarios where strong correlations develop between subsystem pairs, enabling us to study the conditions under which entanglement between A and C is significant or negligible. Specifically, we’re interested in two distinct physical regimes characterized by the relative strength of the self-squeezing parameter $\lambda$ compared to the oscillator frequencies $\omega$:
  1. 1.
    The heavy mediator regime (HMR), where the self-squeezing strength of the mediator oscillator B approaches its natural frequency, i.e., $\lambda_{B}\to\omega_{B}$.
  2. 2.
    The light mediator regime (LMR), where the self-squeezing strengths of the peripheral oscillators A and C approach their natural frequencies $\lambda_{A}\to\omega_{A}$.
As we show explicitly, these two regimes give rise to qualitatively different entanglement and correlation dynamics. The total Hamiltonian (in units of $\hbar$) is
@@
where the superscript ${(0)}$ denote the free Hamiltonians of the individual HOs along with self squeezing interaction strength ($\lambda_{A},\lambda_{B}$), while the last two terms describe the interactions strengths $(g_{A},g_{B})$:
$\hat{H}^{(0)}_{A}$ $=\omega_{A}\left(a^{{\dagger}}a+\mathbb{I}/2\right)+\lambda_{A}\left(a^{2}+a^{{\dagger}2}\right)/2$ (2a)
$\hat{H}^{(0)}_{B}$ $=\omega_{B}\left(b^{{\dagger}}b+\mathbb{I}/2\right)+\lambda_{B}\left(b^{2}+b^{{\dagger}2}\right)/2$ (2b)
$\hat{H}^{(0)}_{C}$ $=\omega_{A}\left(c^{{\dagger}}c+\mathbb{I}/2\right)+\lambda_{A}\left(c^{2}+c^{{\dagger}2}\right)/2$ (2c)
$\hat{H}_{AB}$ $=g_{A}\left(a^{{\dagger}}b+ab^{{\dagger}}\right)+g_{A}\left(ab+a^{{\dagger}}b^{{\dagger}}\right)$ (2d)
$\hat{H}_{BC}$ $=g_{B}\left(b^{{\dagger}}c+bc^{{\dagger}}\right)+g_{B}\left(bc+b^{{\dagger}}c^{{\dagger}}\right)$ (2e)
The annihilation and creation operators ($a,b,c$) and ($a^{{\dagger}},b^{{\dagger}},c^{{\dagger}}$) satisfy the standard commutation relations: $\left[a,a^{{\dagger}}\right]=\mathbb{I};~{}\left[b,b^{{\dagger}}\right]=\mathbb{I};~{}\left[c,c^{{\dagger}}\right]=\mathbb{I}$. To study the system’s dynamics, we employ the covariance matrix formalism [40]. (See Appendix (A) for details.) First, we express Hamiltonian (1) as:
@@
$\bm{U}=\begin{pmatrix}\omega_{A}&g_{A}&0\\ g_{A}&\omega_{B}&g_{B}\\ 0&g_{B}&\omega_{A}\\ \end{pmatrix}\quad\bm{V}=\begin{pmatrix}\lambda_{A}&g_{A}&0\\ g_{A}&\lambda_{B}&g_{B}\\ 0&g_{B}&\lambda_{A}\end{pmatrix}$ (4)
@@
The Hamiltonian in Eq. (1) is quadratic in the Bosonic creation and annihilation operators, guaranteeing that the time evolution preserves Gaussianity [41]. Consequently, if the system is initially in a Gaussian state, it remains Gaussian throughout [42]. We therefore restrict our analysis to Gaussian states ($\hat{\rho}_{g}$), which are fully characterized by their first and second moments of the quadrature operator vector $\hat{\mathbb{X}}$ and are defined as:
$\mathbf{d}_{i}=\langle\hat{\mathbb{X}}_{i}\rangle_{g};~{}~{}\bm{\sigma}_{nm}=\langle\{\hat{\mathbb{X}}_{n},\hat{\mathbb{X}}^{{\dagger}}_{m}\}\rangle{g}-2\langle\hat{\mathbb{X}}_{n}\rangle_{g}\langle\hat{\mathbb{X}}_{m}^{{\dagger}}\rangle_{g}$ (5)
[…]
The time evolution of the moments is governed by the symplectic transformation $S_{H}(t)=\exp(\Omega Ht)$, where the canonical symplectic form for this complex-operator basis is $\Omega=-i\,\text{diag}(1,1,1,-1,-1,-1)$. This transformation preserves the symplectic structure, $S_{H}(t)\,\Omega\,S_{H}^{T}(t)=\Omega$, thereby ensuring that the commutation relations remain invariant. The first and second moments evolve as:
$\mathbf{d}(t)=S_{H}(t)\,\mathbf{d}(0),\quad\bm{\sigma}(t)=S_{H}(t)\,\bm{\sigma}(0)\,S_{H}^{\dagger}(t).$ (6)
Since first moments do not affect quantum correlations such as entanglement or discord, we set $\mathbf{d}(0)=0$ without loss of generality and focus on the evolution of the covariance matrix $\bm{\sigma}(t)$.
@@
To investigate the development of quantum correlations in this system, we study the time evolution of the covariance matrix $\bm{\sigma}(t)$ for the Hamiltonian (1). In realistic experimental scenarios — such as the proposed BMV-type setups — gravitationally mediated entanglement between two massive quantum systems (e.g., A and C) would be inferred from measurements performed over multiple experimental runs. As such experiments access only ensemble-averaged data, we evaluate long-time averages of ($\mathcal{E}_{\mathcal{N}}$) and ($\mathcal{D}$), computed over timescales much longer than the inverse normal-mode frequencies. These long-time averages qualitatively capture the entanglement structure mediated by oscillator B, reflecting what would be observed in ensemble measurements across many experimental runs.
III Time evolution of the Hamiltonian (1)
@@
The time evolution of the covariance matrix $\bm{\sigma}(t)$ is obtained from the system’s normal mode frequencies $(k_{1},k_{2},k_{3})$ and the Bogoliubov coefficients $\bm{\alpha}$ and $\bm{\beta}$, which transforms $H$ into a diagonal form corresponding to the Williamson normal form [43], combined with the initial state’s covariance matrix (see Appendix (B) for details). To analyze quantum correlations between subsystems A and C, we trace out the mediator B from the time-evolved covariance matrix $\bm{\sigma}(t)$ to obtain the reduced state $\bm{\sigma}_{A-C}(t)$. The logarithmic negativity $\mathcal{E}_{\mathcal{N}}(t)$ and Gaussian quantum discord $\mathcal{D}(t)$ are then computed using standard techniques for Gaussian states [24, 25] (see Appendix (C) for details).
@@
where each |$\ket{r,\theta}$ is a single-mode squeezed vacuum state:
|$\ket{r,\theta}=S(\zeta)\ket{0},\quad S(\zeta)=\exp{\left[(\zeta^{*}\hat{a}^{2}-\zeta\hat{a}^{\dagger 2})/2\right]}$ ) (8)
Here, $\zeta=re^{i\theta}$ is the complex squeezing parameter, with $r$ ($\theta$) as the squeezing strength (angle) [44]. While analytical solutions for $\mathcal{E}_{\mathcal{N}}(t)$ and $\mathcal{D}(t)$ are, in principle, possible for general Gaussian initial states, the complexity in asymmetric parameter regimes makes this approach cumbersome. Therefore, our primary method is numerical simulation to explore the full parameter space.
To gain physical insight, we perform a calculation in the weak coupling limit for the simplest configuration: an initial vacuum state ($r_{1}=r_{2}=r_{3}=0,\theta_{1}=\theta_{2}=\theta_{3}=0$), frequency-degenerate oscillators ($\omega_{A}=\omega_{B}=\omega$), symmetric couplings to the mediator ($g_{A}=g_{B}=g$), and no local squeezing interactions ($\lambda_{A}=\lambda_{B}=0$). In this regime, we consider the limit $\lambda_{p}\equiv\sqrt{2}g/\omega\ll 1$ limit. We find that the long-time averaged logarithmic negativity $\langle\mathcal{E}_{\mathcal{N}}\rangle$ between subsystems A and C takes the form:
$\langle\mathcal{E}_{\mathcal{N}}\rangle\approx\frac{\lambda_{p}}{\pi\ln 2}=\frac{\sqrt{2}}{\pi\ln 2}\frac{g}{\omega}$ (9)
(see Appendix (D) for details). Interestingly, in the weak coupling regime, the average entanglement varies inversely with the mediator frequency $\omega$, implying that lower-frequency mediators facilitate stronger effective correlations. However, the apparent divergence as $\omega\to 0$ is unphysical because the weak coupling condition ($\lambda_{p}\ll 1$) requires $g$ to vanish at least proportionally to $\omega$. This behavior is nonetheless reminiscent of entanglement enhancement seen in bipartite systems featuring zero modes [45, 46].

### v3

# Beyond Entanglement: Diagnosing quantum mediator dynamics in gravitationally mediated experiments

@@ Abstract
No experimental test to date has provided conclusive evidence on the quantum nature of gravity. Recent proposals, such as the BMV experiment, suggest that generating entanglement could serve as a direct test. Motivated by these proposals, we study a system of three-harmonic oscillator system, with the mediator oscillator operating in two distinct parameter regimes: a heavy mediator regime and a light mediator regime. These regimes induce qualitatively different entanglement dynamics between the terminal oscillators. Crucially, distinguishing these regimes experimentally remains challenging when relying solely on entanglement measures. We demonstrate that the dynamical fidelity susceptibility offers a viable and sensitive probe to contrast the regimes in practice. Our results provide testable signatures for optomechanical and trapped-ion platforms simulating gravitational interactions, and provide new avenues to characterize quantum-gravity-inspired systems beyond entanglement-based protocols.
@@ I Introduction
The Bose-Marletto-Vedral (BMV) proposal Bose et al. [2017], Marletto and Vedral [2017] has emerged as a promising avenue to experimentally probe the quantum nature of gravity. It proposes that two mesoscopic masses, each prepared in a spatial superposition, would become entangled via their gravitational interaction if and only if gravity is quantum, a signature absent if mediated by a classical field Oppenheim et al. [2023], Bose et al. [2022, 2025]. While conceptually elegant, the BMV protocol faces dual challenges: Experimental challenges include maintaining coherence and detecting extremely weak gravity induced entanglement Castro-Ruiz et al. [2020], Millen and Stickler [2020], Westphal et al. [2021], Carney et al. [2021], Fuchs et al. [2024]. Theoretical ambiguities persist: studies question whether entanglement is an unambiguous signature of quantum gravity, given that classical systems can mimic certain statistical correlations or even Bell-local correlations under certain conditions Coradeschi et al. [2021], Carney et al. [2019], Altamirano et al. [2018], Ma et al. [2022], Belenchia et al. [2022], Fragkos et al. [2022], Hanif et al. [2024], Galley et al. [2023], Martín-Martínez and Perche [2023], Pedernales and Plenio [2023], Etezad-Razavi and Hardy [2024]. Motivated by these crucial interpretational challenges in gravitationally-mediated entanglement experiments, we develop and analyze a theoretical model of three harmonically coupled oscillators, that effectively mimic the BMV setup with generic quadratic couplings. Our investigation delineates what mediator-induced entanglement can genuinely reveal about the fundamental nature of the mediating interaction itself.
In this work, we resolve this ambiguity through a solvable three-coupled harmonic oscillator (HO) model, where oscillator $B$ acts as a tunable mediator. Our construction mirrors the essential assumptions of the BMV proposal: the initial global state is separable, and subsystems $A$ and $C$ couple only locally to the mediator $B$, never directly to each other. As a result, any entanglement observed between $A$ and $C$ can arise only through the mediator. This allows us to test, in a controlled setting, the central BMV idea that mediator-induced entanglement can act as an indicator of the mediator’s quantum nature. While our three-oscillator setup is not a literal model of gravity — for instance, the mediator is a massive oscillator, not a massless graviton — it serves as a powerful analogue framework. The mediator’s mass effectively sets the interaction range, enabling us to explore distinct dynamical regimes and draw physical insights into BMV-type scenarios.
Within this framework, we explore two contrasting regimes: one where the mediator itself is highly squeezed, and another where the outer oscillators are predominantly squeezed. We first derive an analytical expression for the entanglement (quantified by Logarithmic Negativity ($\mathcal{E}_{\mathcal{N}}$)  Plenio [2005], Vidal and Werner [2002]) in the most symmetric weak-coupling scenario. For generic couplings and various initial states, we then employ numerical simulations to precisely map the entanglement patterns generated in these two extreme regimes. This allows us to identify the inherent difficulties in using entanglement as the sole resource to characterize the mediator. Beyond entanglement, we also analyze the behavior of quantum discord ($\mathcal{D}$) and introduce dynamical fidelity susceptibility ($\chi_{F})$ Chandran and Shankaranarayanan [2024] as a novel diagnostic tool. We show that $\chi_{F}$ clearly discriminates between the two regimes. Finally, by prescribing parameter tuning (mediator mass, coupling strengths) and monitoring $\chi_{F}$, we provide practical guidance to avoid null-result pitfalls in BMV-type experiments.
@@ II The model and set up
Consider three 1-D quantum HOs, as depicted in Fig. (1). We treat the oscillators A and C as identical HOs with same shape, mass ($m_{A}$) and frequency ($\omega_{A}$) and the oscillator B acts as a local mediator connecting A and C, which do not interact directly. The HOs A and C and their interaction with the mediator B is modelled via the Hamiltonian (in units of $\hbar$)
@@
where the superscript (0){(0)} denote the free Hamiltonians of the individual HOs along with self squeezing interaction strength ($\lambda_{A},\lambda_{B}$), while the last two terms describe the interactions strengths $(g_{A},g_{B})$:
$\hat{H}^{(0)}_{i}$ $=$ $\omega_{i}\left(a_{i}^{{\dagger}}a_{i}+\mathbb{I}/2\right)+\lambda_{i}\left(a_{i}^{2}+a_{i}^{{\dagger}2}\right)/2\,$, (2a)
$\hat{H}_{ij}$ $=$ $g_{i}\left(a_{i}^{{\dagger}}a_{j}+a_{i}a_{j}^{{\dagger}}\right)+g_{i}\left(a_{i}a_{j}+a_{i}^{{\dagger}}a_{j}^{{\dagger}}\right)\,$, (2b)
where $i,j=\textit{A,B,C}$ and $a_{i}=(a,b,c)$ respectively for $i=A,B,C.$ The annihilation and creation operators ($a,b,c$) and ($a^{{\dagger}},b^{{\dagger}},c^{{\dagger}}$) satisfy the standard bosonic commutation relations. The Hamiltonian incorporates three types of interactions: self-squeezing, parametric squeezing, and mode mixing. The interaction between the mediator and the other HOs resembles those in the Caldeira-Leggett model, coupling subsystems via terms like $x_{A}x_{B}$ or $x_{B}x_{C}$ Caldeira and Leggett [1983]. These inter-oscillator couplings decompose into two qualitatively distinct components: mode mixing, which conserves particle number (e.g., $a^{\dagger}b+ab^{\dagger}$), and parametric squeezing, which does not (e.g., $ab+a^{\dagger}b^{\dagger}$Rugar and Grütter [1991]. The parametric squeezing and mode-mixing terms in our Hamiltonian are reminiscent of the effective Newtonian interaction that would arise between two point masses at large separation. However, unlike a direct two-body coupling that would directly entangle $A$ and $C$, in our construction the interaction is realized locally via the mediator $B$, while still retaining the essential features of a Newtonian potential. This decomposition facilitates non-trivial energy redistribution and building of correlations across subsystems.
The inclusion of self-squeezing terms in the above Hamiltonian, similar to those observed in optomechanical systems, is motivated by the non-linear nature of gravity, where even in the linear limit, phenomena like gravitational memory exist Zel’dovich and Polnarev [1974], Braginskii and Thorne [1987], Chakraborty et al. [2025]. Thus, self-squeezing is crucial for model aiming to shed light on gravitational interactions. Taken together, these ingredients ensure that the model retains the key structural features associated with gravitational mediation: interactions are strictly local, the coupling structure echoes that of a Newtonian potential, and nonlinear self-effects are incorporated through the squeezing terms. Crucially, this controlled setup allows us to probe in a precise way what the subsequent emergence of entanglement between A and C actually signals about the quantum character of the mediator. To study the system’s dynamics, we employ the covariance matrix formalism Eisert et al. [2010] (See Appendix (A) for details.) First, we express Hamiltonian (1) as:
@@
$\bm{U}=\begin{pmatrix}\omega_{A}&g_{A}&0\\ g_{A}&\omega_{B}&g_{B}\\ 0&g_{B}&\omega_{A}\\ \end{pmatrix}\quad\bm{V}=\begin{pmatrix}\lambda_{A}&g_{A}&0\\ g_{A}&\lambda_{B}&g_{B}\\ 0&g_{B}&\lambda_{A}\end{pmatrix}\,$. (4)
@@
The Hamiltonian (1) is quadratic, guaranteeing that the time evolution preserves Gaussianity Adesso et al. [2014]. Consequently, if the system is initially in a Gaussian state, it remains Gaussian throughout Chandran and Shankaranarayanan [2023]. We therefore restrict our analysis to Gaussian states ($\hat{\rho}_{g}$), which are fully characterized by their second moments of the quadrature operator vector $\hat{\mathbb{X}}$ and are defined as:
$\bm{\sigma}_{nm}=\langle\{\hat{\mathbb{X}}_{n},\hat{\mathbb{X}}^{{\dagger}}_{m}\}\rangle_{g}-2\langle\hat{\mathbb{X}}_{n}\rangle_{g}\langle\hat{\mathbb{X}}_{m}^{{\dagger}}\rangle_{g}$ (5)
where $\langle\cdot\rangle_{g}$ denotes the expectation value of the operator with respect to state $\hat{\rho}_{g}$ and $\cdot$} is the anticommutator.
@@
The time evolution of the moments is governed by the symplectic transformation $S_{H}(t)=\exp(\Omega Ht)$, where the canonical symplectic form for this complex-operator basis is $\Omega=-i\,\text{diag}(1,1,1,-1,-1,-1)$. This transformation preserves the symplectic structure, $S_{H}(t)\,\Omega\,S_{H}^{T}(t)=\Omega$, thereby ensuring that the commutation relations remain invariant. The second moments evolve as: $\bm{\sigma}(t)=S_{H}(t)\,\bm{\sigma}(0)\,S_{H}^{\dagger}(t)$. To investigate the development of quantum correlations in this system, we analyze the time evolution of the covariance matrix $\bm{\sigma}(t)$ for the Hamiltonian (1). In realistic experimental scenarios — such as the BMV-type setups — gravitationally mediated entanglement between two massive quantum systems (e.g., A and C) would be inferred from measurements performed over multiple experimental runs. Since experiments access only ensemble-averaged data, we evaluate long-time averages of ($\mathcal{E}_{\mathcal{N}}$), computed over timescales much longer than the inverse normal-mode frequencies. This approach qualitatively captures the entanglement structure mediated by oscillator B, reflecting what would be observed in ensemble measurements across many experimental runs.
III Time evolution of the Hamiltonian
@@
The time evolution of the covariance matrix $\bm{\sigma}(t)$ is obtained from the system’s normal mode frequencies $(k_{1},k_{2},k_{3})$ and the Bogoliubov coefficients $\bm{\alpha}$ and $\bm{\beta}$, combined with the initial state’s covariance matrix (see Appendix (B) for details). To analyze quantum correlations between subsystems A and C, we trace out the mediator B from the time-evolved covariance matrix $\bm{\sigma}(t)$ to obtain the reduced state $\bm{\sigma}_{A-C}(t)$. The logarithmic negativity $\mathcal{E}_{\mathcal{N}}(t)$ and quantum discord $\mathcal{D}(t)$ are then computed using standard techniques for Gaussian states Vidal and Werner [2002].
@@
[…]
Here, $\zeta=re^{i\theta}$ is the complex squeezing parameter, with $r$ ($\theta$) as the squeezing strength (angle) Caves [1981]. While analytical solutions for $\mathcal{E}_{\mathcal{N}}(t)$ and $\mathcal{D}(t)$ are, in principle, possible for general Gaussian initial states, the complexity in asymmetric parameter regimes makes this approach cumbersome. Therefore, our primary method is numerical simulation to explore the full parameter space.
To gain physical insight, we obtain an analytical expression for $\mathcal{E}_{\mathcal{N}}$ in the weak coupling limit under a highly simplified configuration. This specific scenario assumes an initial vacuum state ($r_{1}=r_{2}=r_{3}=0,\theta_{1}=\theta_{2}=\theta_{3}=0$), frequency-degenerate oscillators ($\omega_{A}=\omega_{B}=\omega$), symmetric couplings to the mediator ($g_{A}=g_{B}=g$), and no local squeezing interactions ($\lambda_{A}=\lambda_{B}=0$). In this regime, we consider the limit $\lambda_{p}\equiv\sqrt{2}g/\omega\ll 1$. We find that the long-time averaged logarithmic negativity $\langle\mathcal{E}_{\mathcal{N}}\rangle$ between subsystems A and C takes the form:
$\langle\mathcal{E}_{\mathcal{N}}\rangle\approx{\lambda_{p}}/{(\pi\ln 2)}=(\sqrt{2}g)/({\pi\ln 2\omega})$ (8)
Interestingly, in the weak coupling regime, the average entanglement varies inversely with the mediator frequency $\omega$, implying that lower-frequency mediators facilitate stronger effective correlations. However, the apparent divergence as $\omega\to 0$ is unphysical because the weak coupling condition ($\lambda_{p}\ll 1$) requires $g$ to vanish at least proportionally to $\omega$. This behavior is nonetheless reminiscent of entanglement enhancement seen in bipartite systems featuring zero modes Chandran and Shankaranarayanan [2019], Jain et al. [2021].
Figure 2: Contour plots of the long-time averaged logarithmic negativity $\langle\mathcal{E}_{\mathcal{N}}\rangle$ in the two mediator regimes. Top: (HMR) Bottom: (LMR). Parameter choices and initial squeezed states are specified in the text.